OpenStudy (anonymous):

dy/dx = 2y^2 and if y = -1 when x = 1. then when x = 2, y=?

OpenStudy (anonymous):

dy/y^2 = 2*dx, now integratin both sides -1/y = 2x + C when x = 1, y=-1 -1/-1 = 2*1 + C 1 = 2 + C C = -1, so... -1/y = 2x -1 When x=2: -1/y = 4 -1 -1/y = 3 y = -1/3

OpenStudy (solomonzelman):

but there is no x in the derivative....

OpenStudy (solomonzelman):

oh, integration on both sides, didn't think of that

OpenStudy (solomonzelman):

yeah, I will make sure to include this type of problems as well, ty

OpenStudy (anonymous):

:)

OpenStudy (solomonzelman):

I thought it was one of those, when the differential dx=... and f(x)=... find dy,. Looked at it wrong.....

OpenStudy (kamibug):

\(\huge\bf\color{seagreen}{Welcome~to~Openstudy!~:)}\) @audiq7

OpenStudy (solomonzelman):

teal